# On polynomial automorphisms of Nagata type

### v1

@@ Abstract.
We define a family of polynomial ring homomorphisms generalising the Nagata automorphism, and relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation. We also give a condition on that family to get wild automorphisms.
@@
polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE
@@ 1. Introduction
In what follows $F$ will be a field of characteristic 00 unless otherwise stated. The polynomials:
$f(x,y,z)$ $=x-2y(zx+y^{2})-z(zx+y^{2})^{2},$
$g(x,y,z)$ $=y+z(zx+y^{2}),$
$h(x,y,z)$ $=z.$
define an automorphism on $F[x,y,z]$ called Nagata automorphism [1]. The determinant of the Jacobian matrix of this automorphism es $1$. In the following paragraph we recall the notion of tame automorphisms of polynomials. Let $1\leq i\leq n$, the $i$th elementary automorphism $\Phi^{(i)}\colon F^{n}\to F^{n}$ is the automorphism defined as
$\Phi^{(i)}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}$
and its inverse is the automorphism
${\Phi^{(i)}}^{-1}(x_{1},\dots,x_{n})=\Big{(}x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big{)}\,.$
One may notice that the Jacobian matriz of $\Phi^{(i)}$ has determinant $1$. An automorphism is tame if it is a finite composition of elementary and linear automorphisms. An automorphism that is not tame is called wild. In 1972, Nagata conjectured that the Nagata automorphism is wild. In [2, Corollary 9], the authors proved this conjecture. In this work, we define ring homomorphisms on $F[x,y,z]$, given in (2), extending the Nagata automorphism which are wild automorphisms under some assumptions, see Corollary 1). We also relate the condition of non-vanishing of their Jacobian determinant with a certain linear partial differential equation, see Theorem 1. As a consequence, we prove that those ring homomorphisms on $F[x,y,z]$ with nonzero Jacobian determinant, are in fact automorphisms, see Theorem 2.
@@ 2. Nagata polynomial map
The triplet $(f,g,h)$ induces a ring homomorphism on $F[x,y,z]$. Note that $(f,g,h)$ is the identity when one takes $\varphi=0$, so $(f,g,h)$ may be seen as a perturbation of the identity. If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism. The Milnor number of $(f,g,h)$ is the dimension
@@
If $(f,g,h)$ is a ring homomorphism as in (2), then $\mu(f,g,h)=1$.
@@
Note that $(f,g,h)$ is not always an automorphism for any $\varphi$ despite the fact that $\mu(f,g,h)=1$ by Lemma 1. For instance, if $\varphi=x$ or $\varphi=y$, then $(f,g,h)$ is not an automorphism. In the first case, $(f,g,h)$ sends both points $(0,0,1)$ and $(-1,1,1)$ to $(0,0,1)$. Theorem 2 provides a necessary and sufficient condition on $\varphi$ such that $(f,g,h)$ is an automorphism.
@@
The Jacobian matrix $J(f,g,h)$ has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear PDE:
$-2y\varphi_{x}+z\varphi_{y}=0\,.$
@@
$f_{y}$ $=-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y},$
@@
$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
@@
$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi_{x}-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
If $-2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
@@
However this equality is only possible when $c=1$, therefore $-2y\varphi_{x}+z\varphi_{y}=0$. ∎
@@
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the ring homomorphism $(f,g,h)$ given in (2) is an automorphism on $F[x,y,z]$ with inverse:
@@
For a polynomial $\psi$, we denote by $\overline{\psi}$ the highest degree homogeneous component of $\psi$.
@@
Let $\varphi\in F[x,y,z]$ and suppose that $(f,g,h)$ is an automorphism on $F[x,y,z]$ of the form (2). If one the partial derivatives $\overline{\varphi}_{x},\overline{\varphi}_{y}$ is nonzero, then the automorphism $(f,g,h)$ is wild.
@@
By [3, Corollary 2] or [2, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent. On the other hand, from the hypothesis $\overline{\varphi}\in F[x,y,z]$ is a homogeneous polynomial of positive degree such that $\overline{\varphi}\not\in F[z]$, it is not difficult to see that none of the polynomials $\overline{f},\overline{g},\overline{h}$ belongs to the subalgebra generated by the other two polynomials. Therefore, by [2, Corollary 8] $(f,g,h)$ is wild. ∎
Definition 1.
Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$. For $k\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$. A polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
  1. (1)
    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
  2. (2)
    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$. The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$. We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(p)$. In particular, if $v=(1,\dots,1)$ then $\deg(h)=d_{v}(h)$ and $h^{v}=\overline{h}$.
  3. (3)
    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
@@
(3) $\mathrm{supp}(\varphi)=\Big{\{}(k_{1}-n,2n,k_{1}+k_{2}-n):k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big{\}}\,.$
@@
$\deg(\varphi)$ $=\max\Big{\{}k_{1}-n+2n+k_{1}+k_{2}-n:(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big{\}}$
$=\max\Big{\{}2k_{1}+k_{2}:(k_{1},k_{2})\in\mathrm{supp}(p)\Big{\}}$
@@
If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}(x,y,z)=\overline{p}(xz+y^{2},z)$. Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$, then by a immediate computation one has
@@
However, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Moreover, by direct computation one has
$\overline{\varphi}=\left(xz+y^{2}\right)^{2}+\left(xz+y^{2}\right)z^{2}=p^{v}(xz+y^{2},z)\,.$
@@
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, then the corresponding ring homomorphism $(f,g,h)$ given in (2) is a wild automorphism.
@@
Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[s,t]$ as in Example 1. Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding ring homomorphism $(f,g,h)$ is a wild automorphism.
@@ 3. Statement of the main results
The linear PDE
(5) $-2y\varphi_{x}+z\varphi_{y}=0$
has as general solution in $F[x,y,z]$, polynomials that can be written as:
$\varphi=p(xz+y^{2},z)\,,$
where $p\in F[t_{1},t_{2}]$.
Theorem 2.
Let $\varphi\in F[x,y,z]$ and set
$f(x,y,z)$ $=x-2y\varphi-z\varphi^{2}\,,$
$g(x,y,z)$ $=y+z\varphi\,,$
$h(x,y,z)$ $=z\,.$
Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism on $F[x,y,z]$. In this case, there exists a polynomial $p\in F[t_{1},t_{2}]$ such that $\varphi=p(xz+y^{2},z)$, moreover the inverse of $(f,g,h)$ is given in (2).
The proofs of these results are given after the following previous results.
@@
Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear PDE:
(6) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,x)\varphi_{z}=0\,,$
where each $\varphi^{k}\in F[x,y.z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree. Then each homogeneous component $\varphi^{k}$ is also a solution of the equation (6).
@@
$\sum_{k=0}^{n}\Big{(}A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big{)}=0$
which is a sum of homogeneous polynomials of different degrees, therefore:
$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,.$
By the previous lemma, to find the general solution of the equation (5), it is enough to deal with its homogeneous polynomials components. Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ is written as:
(7) $\varphi=\varphi^{n}+\varphi^{n-1}z+\varphi^{n-2}z^{2}+\dots+\varphi^{1}z^{n-1}+\varphi^{0}z^{n}=\sum_{k=0}^{n-1}\varphi^{n-k}z^{k}+\varphi^{0}z^{n}$
with $\varphi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
@@
Let $\varphi$ be a homogeneous polynomial in $F[x,y,z]$ of degree $n$ that is a solution of the linear PDE (5). Then, $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
$2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}$
@@
$\varphi_{x}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{x}z^{k}=\varphi_{x}^{n}+\sum_{k=1}^{n-1}\varphi^{n-k}_{x}z^{k}\,,$
$\varphi_{y}$ $=\sum_{k=0}^{n-1}\varphi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\varphi^{n-k}_{y}z^{k}+\varphi^{1}_{y}z^{n-1}\,.$
@@
$-2y\varphi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\varphi^{n-k-1}_{x}+\varphi^{n-k}_{y}\right)z^{k+1}+\varphi^{1}_{y}z^{n}=0\,.$
This equation gives the equalities $\varphi^{n}_{x}=0$ and $\varphi^{1}_{y}=0$, so $\varphi^{n}=ay^{n}$ and $\varphi^{1}=bx$ for some constants $a,b\in F$. Moreover,
(8) $2y\varphi^{n-k-1}_{x}=\varphi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
@@
    If $\varphi$ has degree $2n$, then
    $\varphi=a(xz+y^{2})^{n}+z\psi$
    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (5) of degree $12n-$ 1.
@@
    If $\varphi$ has degree $12n+$ 1, then
    $\varphi=z\psi$

### v2

@@ Abstract.
We define a family of polynomial ring homomorphisms generalizing the well-known Nagata automorphism. We establish necessary and sufficient conditions under which these homomorphisms are automorphisms, and verify that they satisfy the Jacobian conjecture. Additionally, we provide a necessary condition within this family to obtain wild automorphisms, and independently derive a property related to the upper semicontinuity of the Łojasiewicz exponent at infinity.
@@
polynomial automorphisms, Nagata automorphism, wild automorphisms, linear PDE, Łojasiewicz exponent
@@ 1. Introduction
In what follows $F$ will be a field of characteristic 0 unless otherwise stated. The polynomials
(1) $\begin{split}f(x,y,z)&=x-2y(zx+y^{2})-z(zx+y^{2})^{2},\\ g(x,y,z)&=y+z(zx+y^{2}),\\ h(x,y,z)&=z.\end{split}$
define an automorphism of $F[x,y,z]$ called Nagata automorphism [7]. The determinant of the Jacobian matrix of this automorphism turns out to be $1$.
An automorphism is tame if it is a finite composition of elementary and linear automorphisms. An automorphism that is not tame is called wild [12].
In 1972, Nagata conjectured that the aforementioned automorphism is wild. This conjecture was proved in [12, Corollary 9].
In this work, we define a family of ring homomorphisms on $F[x,y,z]$, given in (2), generalizing the Nagata automorphism, which includes the family of automorphisms constructed by Bass [1] in some sense, but is quite different to the family of automorphisms constructed in [6, (1)]. Our family provides wild automorphisms under some assumptions, see Corollary 1. In Theorem 1, we solve a linear partial differential equation, whose solution is related to the non-vanishing of the Jacobian determinant of those automorphisms. As a consequence, we obtain a necessary and sufficient condition on homomorphisms of Nagata type with nonzero Jacobian determinant to be automorphisms, which is a particular case of the famous Jacobian conjecture, see Corollary 3.
Acknowledgments. We are grateful to Carles Bivià-Ausina for valuable comments on an early version of this article.
@@ 2. Nagata polynomial map
In the following paragraph we recall the notion of tame automorphisms of polynomials.
Fix a natural number $n\geq 2$. For each index $1\leq i\leq n$, the $i$th elementary automorphism of $F^{n}$ on itself is the automorphism defined as
$(x_{1},\dots,x_{n})\mapsto\Big(x_{1},\dots,x_{i-1},x_{i}+f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big)\,$,
where $f$ is a polynomial map in n1n-1 variables, and its inverse is the automorphism
$(x_{1},\dots,x_{n})\mapsto\Big(x_{1},\dots,x_{i-1},x_{i}-f(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n}),x_{i+1},\dots,x_{n}\Big)\,.$
It turns out that the Jacobian matrix of each elementary automorphism has determinant $1$.
Since $F$ is an infinite field, there is a one-to-one correspondence between polynomial maps $F^{n}\to F^{n}$ and $n$-tuples of polynomials in $F[x_{1},\dots,x_{n}]$. In what follows, we shall focus on the case $n=3$.
Definition 1.
@@
The triplet $(f,g,h)$ induces an ring homomorphism on $F[x,y,z]$ which will be called Nagata homomorphism.
Note that $(f,g,h)$ is the identity when one takes $\varphi=0$, so $(f,g,h)$ may be seen as a perturbation of the identity. If $\varphi=xz+y^{2}$ then $(f,g,h)$ is the Nagata automorphism (1). If $\varphi=\lambda(xz+y^{2})$ with $\lambda\in F$, then we obtain a family of automorphisms similar to those constructed by Bass [1], [5, 3.8.1].
The Milnor number of $(f,g,h)$ is the dimension
@@
If $(f,g,h)$ is a Nagata homomorphism, then $\mu(f,g,h)=1$.
@@
Note that the Nagata homomorphism $(f,g,h)$ is not always an automorphism for any $\varphi$, despite the fact that $\mu(f,g,h)=1$ by Lemma 1. For instance, if $\varphi=x$, then $(f,g,h)$ is not an automorphism, because as a map $(f,g,h)$ sends both points $(0,0,1)$ and (1,1,1)(-1,1,1) to $(0,0,1)$. Similarly, $(f,g,h)$ is not an automorphism for $\varphi=y$. In contrast, Corollary 3 provides a necessary and sufficient condition on $\varphi$, in order that its corresponding Nagata homomorphism $(f,g,h)$ is an automorphism.
@@
The Jacobian matrix $J(f,g,h)$ of the Nagata homomorphism $(f,g,h)$, as in (2), has nonzero determinant in $F$ if and only if $\varphi$ satisfies the linear partial differential equation:
$2y\varphi_{x}+z\varphi_{y}=0\,.$
@@
$f_{y}$ $=-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y},$
@@
$\begin{pmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y}&f_{z}\\ z\varphi_{x}&1+z\varphi_{y}&g_{z}\\ 0&0&1\end{pmatrix}$
@@
$\begin{vmatrix}1-2y\varphi_{x}-2z\varphi\varphi_{x}&-2\varphi-2y\varphi_{y}-2z\varphi\varphi_{y}\\ z\varphi_{x}&1+z\varphi_{y}\end{vmatrix}=1-2y\varphi_{x}+z\varphi_{y}\,.$
If $2y\varphi_{x}+z\varphi_{y}=0$ then the determinant of $J(f,g,h)$ is $1$. Conversely, if the determinant of $J(f,g,h)$ is a constant $c\in F-\{0\}$, then
@@
However this equality is only possible when $c=1$, therefore $2y\varphi_{x}+z\varphi_{y}=0$. ∎
@@
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. Then the corresponding Nagata homomorphism $(f,g,h)$ is an automorphism of $F[x,y,z]$ with inverse:
@@
Definition 2.
Let us fix algebraically independent indeterminates $t_{1},\dots,t_{n}$ over $F$. For $k=(t_{1},\dots,t_{n})\in\mathbb{Z}^{n}_{\geq 1}$, we write $t^{k}$ to denote the monomial $t_{1}^{k_{1}}\cdots t_{n}^{k_{n}}$. Any polynomial $h\in F[t_{1},\dots,t_{n}]$ can be written as $h=\sum_{k}a_{k}t^{k}$.
  1. (1)
    The support of $h$, denoted by $\mathrm{supp}(h)$, is defined as the set
    $\mathrm{supp}(h)=\{k\in\mathbb{Z}^{n}_{\geq 0}:a_{k}\neq 0\}\,.$
  2. (2)
    Let $v=(v_{1},\dots,v_{n})\in\mathbb{Z}^{n}_{\geq 1}$. The degree of $h$ with respect to $v$, denoted by $d_{v}(h)$, is the maximum of the scalar product $\langle v,k\rangle$ such that $a_{k}\neq 0$. We denote by $h^{v}$ the sum of those terms $a_{k}t^{k}$ such that $\langle v,k\rangle=d_{v}(h)$. In particular, for $v=(1,\dots,1)$, we write
    $\overline{h}=h^{v}\,$,
    that is the highest degree homogeneous component of $h$, and in this case we have $d_{v}(h)=\deg(h)$.
  3. (3)
    We say that $h$ is weighted homogeneous with respect to $v$ when $h^{v}=h$.
@@
Let $\varphi\in F[x,y,z]$ and suppose that its corresponding Nagata homomorphism $(f,g,h)$ is an automorphism of $F[x,y,z]$. If one the partial derivatives $\overline{\varphi}_{x},\overline{\varphi}_{y}$ is nonzero, then the automorphism $(f,g,h)$ is wild.
@@
By [13, Corollary 2] or [12, Lemma 1], we deduce that $\overline{f},\overline{g},\overline{h}$ are mutually algebraically independent. On the other hand, from the hypothesis $\overline{\varphi}\in F[x,y,z]$ is a homogeneous polynomial of positive degree such that $\overline{\varphi}\not\in F[z]$, it is not difficult to see that none of the polynomials $\overline{f},\overline{g},\overline{h}$ belongs to the subalgebra generated by the other two polynomials. Therefore, by [12, Corollary 8] $(f,g,h)$ is wild. ∎
@@
(4) $\mathrm{supp}(\varphi)=\Big\{(k_{1}-n,2n,k_{1}+k_{2}-n):\quad k\in\mathrm{supp}(p)\,\textnormal{and}\,\,0\leq n\leq k_{1}\Big\}\,.$
@@
$\deg(\varphi)$ $=\max\Big\{k_{1}-n+2n+k_{1}+k_{2}-n:\quad(k_{1},k_{2})\in\mathrm{supp}(p)\,\textnormal{and}\,0\leq n\leq k_{1}\Big\}$
$=\max\Big\{2k_{1}+k_{2}:\quad(k_{1},k_{2})\in\mathrm{supp}(p)\Big\}$
@@
If $\varphi(x,y,z)=p(xz+y^{2},z)$, then it is not always true that $\overline{\varphi}$ coincides with $\overline{p}(xz+y^{2},z)$. Indeed, consider $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[t_{1},t_{2}]$, hence $\overline{p}=-t_{2}^{3}+t_{1}t_{2}^{2}$, and by an immediate computation one has
@@
On the other hand, it is easy to verify that if $v=(2,1)$, then $d_{v}(p)=4$ and $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$. Finally, by a direct computation one has
$\overline{\varphi}=p^{v}(xz+y^{2},z)\,.$
@@
Let $p\in F[t_{1},t_{2}]$ and set $\varphi=p(xz+y^{2},z)$ in $F[x,y,z]$. If the partial derivative $(p^{v})_{t_{1}}$ is nonzero, where $v=(2,1)$, then its corresponding Nagata homomorphism $(f,g,h)$ is a wild automorphism.
@@
Let $\varphi(x,y,z)=p(xz+y^{2},z)$, where $p(t_{1},t_{2})=t_{1}^{2}-t_{2}^{3}+t_{1}t_{2}^{2}\in F[t_{1},t_{2}]$ as in Example 1. Since $p^{v}=t_{1}^{2}+t_{1}t_{2}^{2}$ and $(p^{v})_{t_{1}}=2t_{1}+t_{1}^{2}\neq 0$, Corollary 1 implies that corresponding Nagata homomorphism $(f,g,h)$ is a wild automorphism.
Łojasiewicz exponent.— Let $K$ be a normed field. For $z=(z_{1},z_{2},\dots,z_{m})\in K^{m}$, we write $|z|=\max_{1\leq i\leq m}|z_{i}|$. If $F=(F_{1},F_{2},\dots,F_{n})\colon K^{m}\to K^{n}$ is a polynomial map, then we set $\deg F=\max_{1\leq i\leq n}\deg F_{i}$. The Łojasiewicz exponent at infinity of a polynomial map $F\colon K^{m}\to K^{n}$, denoted by $\mathcal{L}_{\infty}(F)$, is defined as the supremum of the real numbers $\alpha$ such that there exist two positive real numbers $C$ and $R$ so that
$|F(z)|\geq C|z|^{\alpha}$
for all $|z|\geq R$ in $K^{m}$. One can follow the argument provided in [11, (1.1)] to deduce that
(6) $\mathcal{L}_{\infty}(F)=\dfrac{1}{\deg(F^{-1})}$
for every polynomial automorphism $F$.
For $K=\mathbb{R}$ or $K=\mathbb{C}$, it is well known that $\mathcal{L}_{\infty}(F)$ exists if and only if $F^{-1}(0)$ is compact, and in this case, $\mathcal{L}_{\infty}(F)$ is a rational number [8, Proposition 2.6]. The computation of this number is generally considered a highly nontrivial problem [4, 9, 11, 2, 3].
A polynomial deformation of a polynomial $p(t_{1},\dots,t_{m})\in K[t_{1},\dots,t_{m}]$ is a polynomial $H(t_{1},\dots,t_{m},s_{1},\dots,s_{n})\in K[t_{1},\dots,t_{m},s_{1},\dots,s_{n}]$ such that
$H(t_{1},\dots,t_{m},0,\dots,0)=p(t_{1},\dots,t_{m})\,.$
For simplicity, we write $s=(s_{1},\dots,s_{n})$, so that the polynomial ring $K[t_{1},\dots,t_{m},s_{1},\dots,s_{n}]$ shall be shortly written as $K[t_{1},\dots,t_{m},s]$, and this allows to set
$p_{s}(t_{1},\dots,t_{m})=H(t_{1},\dots,t_{m},s_{1},\dots,s_{n})\,.$
The following is motivated by the lower semicontinuous property of the Łojasiewicz exponent proven for finite holomorphic germs with constant Milnor number [10, Theorem 2.1]. However, for polynomial maps, we expect the upper semicontinuous property under certain conditions, in view of the fact that Nagata homomorphisms have a constant Milnor number, see Lemma 1.
Corollary 2.
Let $p_{s}(t_{1},t_{2})\in K[t_{1},t_{2},s]$ be a polynomial deformation of a polynomial $p(t_{1},t_{2})\in K[t_{1},t_{2}]$. Let $\varphi=p(xz+y^{2},z)$, $\varphi_{s}=p_{s}(xz+y^{2},z)$, and suppose that $(f,g,h)$ and $(f_{s},g_{s},h_{s})$ are their corresponding Nagata automorphism. Then
$\mathcal{L}_{\infty}(f_{s},g_{s},h_{s})\leq\mathcal{L}_{\infty}(f,g,h)$
Proof.
Since $\mathrm{supp}(p)\subseteq\mathrm{supp}(p_{s})$ with respect to the variables $t_{1}$ and $t_{2}$, we have $\deg(p)\leq\deg(p_{s})$, and for $v=(2,1)$ one has
$\deg\varphi=d_{v}(p)\leq d_{v}(p_{s})=\deg\varphi_{s}\,.$
Hence, in view of (6) we have
$\mathcal{L}_{\infty}(f_{s},g_{s},h_{s})$ $=\dfrac{1}{\deg((f_{s},g_{s},h_{s})^{-1})}$
$=\dfrac{1}{2\deg\varphi_{s}+1}$
$\leq\dfrac{1}{2\deg\varphi+1}$
$=\dfrac{1}{\deg((f,g,h)^{-1})}$
$=\mathcal{L}_{\infty}(f,g,h)\,.$
@@ 3. Statement of the main results
The general solution in $F[x,y,z]$ of the linear partial differential equation:
(7) $2y\varphi_{x}+z\varphi_{y}=0$
has the form $\varphi=p(xz+y^{2},z)$, where $p\in F[t_{1},t_{2}]$.
Corollary 3.
Let $\varphi\in F[x,y,z]$ and let $(f,g,h)$ be its corresponding Nagata homomorphism as in (2). Then, the determinant of the Jacobian matrix $J(f,g,h)$ is a nonzero constant in $F$ if and only if $(f,g,h)$ is an automorphism of $F[x,y,z]$. In this case, there exists a polynomial $p\in F[t_{1},t_{2}]$ such that $\varphi=p(xz+y^{2},z)$, moreover the inverse of $(f,g,h)$ is given in (3).
The proofs of these results are given in Section 4. First of all, we shall provide some preliminary results in the next paragraphs.
@@
Let $\varphi=\varphi^{0}+\varphi^{1}+\cdots+\varphi^{n}$ be a solution of the linear partial differential equation:
(8) $A(x,y,z)\varphi_{x}+B(x,y,z)\varphi_{y}+C(x,y,z)\varphi_{z}=0\,$,
where each $\varphi^{k}\in F[x,y,z]$ is the homogeneous component of $\varphi$ of degree $k$ for $0\leq k\leq n$, and $A,B,C$ are homogeneous polynomials of the same degree. Then each homogeneous component $\varphi^{k}$ is also a solution of the equation (8).
@@
$\sum_{k=0}^{n}\Big(A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}\Big)=0\,$,
which is a sum of homogeneous polynomials of different degrees. Therefore,
$A(x,y,z)\varphi^{k}_{x}+B(x,y,z)\varphi^{k}_{y}+C(x,y,x)\varphi^{k}_{z}=0\,\,\,\textnormal{for all}\,\,k=0,1,\dots,n\,$,
as wanted. ∎
By the previous lemma, to find the general solution of the equation (7), it is enough to deal with its homogeneous polynomials components.
Let $\varphi\in F[x,y,z]$ be a homogeneous polynomial of degree $n$, if we consider the canonical isomorphism of rings $F[x,y,z]\cong F[x,y][z]$, then we can suppose that $\varphi$ has the form:
(9) $\varphi=\phi^{n}+\phi^{n-1}z+\phi^{n-2}z^{2}+\dots+\phi^{1}z^{n-1}+\phi^{0}z^{n}=\sum_{k=0}^{n-1}\phi^{n-k}z^{k}+\phi^{0}z^{n}\,$,
with $\phi^{k}\in F[x,y]$ homogeneous polynomials of degree $k$ for $k=0,1,\dots,n$.
@@
Let $\varphi$ be a homogeneous polynomial in $F[x,y,z]$ of degree $n$ that is a solution of the linear PDE (7). Then, $\phi^{n}=ay^{n}$ and $\phi^{1}=bx$, where $a$ and $b$ are constants in $F$, and moreover
$2y\phi^{n-k-1}_{x}=\phi^{n-k}_{y}$
@@
$\varphi_{x}$ $=\sum_{k=0}^{n-1}\phi^{n-k}_{x}z^{k}=\phi_{x}^{n}+\sum_{k=1}^{n-1}\phi^{n-k}_{x}z^{k}\,$,
$\varphi_{y}$ $=\sum_{k=0}^{n-1}\phi^{n-k}_{y}z^{k}=\sum_{k=0}^{n-2}\phi^{n-k}_{y}z^{k}+\phi^{1}_{y}z^{n-1}\,.$
@@
$-2y\phi_{x}^{n}+\sum_{k=0}^{n-2}\left(-2y\phi^{n-k-1}_{x}+\phi^{n-k}_{y}\right)z^{k+1}+\phi^{1}_{y}z^{n}=0\,.$
This equation gives the equalities $\phi^{n}_{x}=0$ and $\phi^{1}_{y}=0$, so $\phi^{n}=ay^{n}$ and $\phi^{1}=bx$ for some constants $a,b\in F$. Moreover,
(10) $2y\phi^{n-k-1}_{x}=\phi^{n-k}_{y}\,\,\,\textnormal{para todo}\,\,\,k=0,1,\dots,n-2\,.$
@@
    If $\varphi$ has degree $2n$, then it has the form:
    $\varphi=a(xz+y^{2})^{n}+z\psi\,$,
    where $a\in F$ and $\psi\in F[x,y,z]$ is a homogeneous polynomial which is also a solution of the equation (7) of degree 2n12n-1.
@@
    If $\varphi$ has degree $2n+1$, then it has the form:
    $\varphi=z\psi\,$,